### v1

@@ Abstract
3D referring expression comprehension (3DREC) and segmentation (3DRES) have overlapping objectives, indicating their potential for collaboration. However, existing collaborative approaches predominantly depend on the results of one task to make predictions for the other, limiting effective collaboration. We argue that employing separate branches for 3DREC and 3DRES tasks enhances the model’s capacity to learn specific information for each task, enabling them to acquire complementary knowledge. Thus, we propose the MCLN framework, which includes independent branches for 3DREC and 3DRES tasks. This enables dedicated exploration of each task and effective coordination between the branches. Furthermore, to facilitate mutual reinforcement between these branches, we introduce a Relative Superpoint Aggregation (RSA) module and an Adaptive Soft Alignment (ASA) module. These modules significantly contribute to the precise alignment of prediction results from the two branches, directing the module to allocate increased attention to key positions. Comprehensive experimental evaluation demonstrates that our proposed method achieves state-of-the-art performance on both the 3DREC and 3DRES tasks, with an increase of 3.27% in Acc@0.5 for 3DREC and 5.22% in mIOU for 3DRES. Our code is available at https://github.com/qzp2018/MCLN.
@@
Figure 1: (a) TGNN predicts the instance mask and obtains the bounding box using the maximum and minimum XYZ values within the segmentation mask. (b) Our method introduces separate branches for 3DREC and 3DRES tasks, along with a Relative Superpoint Aggregation (RSA) module and an Adaptive Soft Alignment (ASA) module to facilitate collaborative training of the two branches.
@@ 1 Introduction
Hence, it is intuitive to explore the feasibility of incorporating distinct branches dedicated to predicting 3DREC and 3DRES outcomes while training them collaboratively. To solve this issue, we use superpoints to facilitate the joint training of 3DREC and 3DRES tasks. The Relative Superpoint Aggregation (RSA) module is proposed to effectively exploit the spatial relationships between superpoints and their surrounding points during the feature aggregation process. Such an approach could potentially mitigate individual shortcomings through mutual compensation. However, merely training these branches together is insufficient, as the presence of separate branches introduces the risk of prediction conflicts, which could jeopardize the overall model performance. Therefore, ensuring alignment in the generated results from both branches becomes crucial. Furthermore, achieving alignment between the branches is vital in facilitating the exchange of bidirectional information between 3DREC and 3DRES branches. This collaborative learning approach empowers the branches to extract valuable information from one another, refining their positioning and contributing to a more effective performance.
@@
  • Our proposed method has achieved state-of-the-art performance in both 3DREC and 3DRES tasks, showcasing a notable improvement of 3.27% Acc@0.5 for 3DREC and a substantial increase of 5.22% mIOU for 3DRES.
@@ 2 Related Work
3D vision and language play pivotal roles in human comprehension of the world, and they serve as significant areas of research for advancing the development of machines with human-like capabilities. Consequently, noteworthy research endeavours in vision and language [42, 19, 20, 21, 31, 32, 30] have surged. Among these, 3D Visual Grounding (3DVG) [46, 44, 41, 17, 7, 5, 24, 9] stands out as a prime exemplar, focusing on the task of precisely localizing objects referred to in natural language descriptions. In a comprehensive scope, 3DVG comprises two fundamental tasks: 3D Referring Expression Comprehension (3DREC) [46, 44, 41, 7, 5, 45, 9] and 3D Referring Expression Segmentation (3DRES) [17, 40, 36]. For 3DREC, most of the existing works [7, 44, 45, 5] take a two-stage approach. Initially, candidate proposals are obtained using a pre-trained 3D object detector or ground truth information. Subsequently, visual and text features are extracted for multi-modal feature fusion, and then the most suitable proposal is selected based on the fused features. Recently, single-stage methods [41, 18] have also shown superior performance, presenting another promising paradigm. Conversely, although the 3DREC task has received considerable attention, the 3DRES task was initially introduced by TGNN [17], with limited exploration in subsequent years. Recently there has been a flurry of work to investigate this area. For example, 3D-STMN [40] employs superpoints for the 3DRES task, greatly improving performance. However, the collaborative work between 3DREC and 3DRES is still very few.
@@ 2.2 Multi-task Learning
Multi-task learning (MTL) [28, 6, 22, 34, 16] serves as a valuable strategy for enhancing task performance, effectively leveraging training data while facilitating the exchange of information among related tasks. MTL has found extensive applications across a diverse range of computer vision tasks. In the 2D vision, many characteristics are shared between detection and segmentation, which have prompted several works [15, 4, 12] to address both tasks within a unified framework. Similarly, REC and RES can also promote each other in joint training [28, 16, 22]. However, due to the limited exploration of 3DRES in recent years, collaborative initiatives between 3DRES and 3DREC have mainly remained uncharted.
TGNN [17] conducts the first attempt to employ a single model to tackle both 3DREC and 3DRES tasks. However, it doesn’t set up a separate branch for the 3DREC task, and this design exhibits several limitations. For example, it ties the performance ceiling for 3DREC solely to 3DRES, overlooking the distinctions between these tasks. Consequently, it severely restricts the interaction between the two tasks and leads to sub-optimal performance. To deal with the above problems, we designed two separate branches for 3DREC and 3DRES and a co-alignment module for joint training.
@@ 3 Method
Figure 2: An overview of our proposed network. Our proposed network comprises a cross-modal encoder, separate decoders for 3DREC and 3DRES, a “Relative Superpoint Aggregation” (RSA) module for superpoint feature generation, and an “Adaptive Soft Alignment” (ASA) module for joint training. The RSA module utilizes relative position relationships for feature aggregation, while the ASA module includes “Soft Alignment” with point-level and mask-level weight adjustments for Dice and Focal losses, and “Adaptive Weight” for dynamic fusion of predicted masks from two branches.
@@ 3.3 Adaptive Soft Alignment
The query mask $\hat{M}_{q}$ acts as an indirect representation of the predicted bounding box. It serves as a connecting link, facilitating the alignment between the bounding box from the 3DREC branch and the mask from the 3DRES branch. This soft alignment scheme enhances the coherence between the two branches. Through the alignment of superpoint masks ${M}_{p}$ and $\hat{M}_{q}$, the indirect preservation of spatial coherence between the outcomes of the 3DREC and 3DRES tasks is effectively achieved. Usually, a Focal loss [25] and a Dice loss [33] will be applied to conduct the alignment between masks:
$L_{\text{Focal}}(M_{p},\hat{M}_{q})=\sum_{i=1}^{m}-\left(\hat{M}_{q}^{i}\cdot(1-M_{p}^{i})^{\gamma}\cdot\log(M_{p}^{i})\right.$ (10)
$\left.+(1-\hat{M}_{q}^{i})\cdot(M_{p}^{i})^{\gamma}\cdot\log(1-M_{p}^{i})\right),$
$L_{\text{Dice}}(M_{p},\hat{M}_{q})=\sum_{i=1}^{m}\left(1-\frac{2|M_{p}^{i}\cap\hat{M}_{q}^{i}|}{|M_{p}^{i}|+|\hat{M}_{q}^{i}|}\right),$ (11)
where $m$ represents the number of superpoints, $M_{p}^{i}$ and $\hat{M}_{q}^{i}$ represents the $i$-th value of $M_{p}$ and $\hat{M}_{q}$, and $\gamma=2$.
@@ 3.4 Overall Loss
$L_{\text{res}}=\beta_{1}(L_{\text{Focal}}^{1}+L_{\text{Focal}}^{2}+L_{\text{Focal}}^{adp}+L_{\text{Focal}}^{point})+\beta_{2}(L_{\text{Dice}}^{1}+L_{\text{Dice}}^{2}+L_{\text{Dice}}^{adp}+L_{\text{Dice}}^{mask}),$ (20)
where $L_{\text{Focal}}^{1},L_{\text{Dice}}^{1}$ denotes the losses associated with the superpoint mask ${M}_{f}$ and the ground truth. Similarly, $L_{\text{Focal}}^{2},L_{\text{Dice}}^{2}$ represents the losses for the superpoint mask ${M}_{q}$. Additionally, $L_{\text{Focal}}^{adp},L_{\text{Dice}}^{adp}$ correspond to the losses for the final predicted superpoint mask ${M}_{f}$. Finally, $L_{\text{Focal}}^{point}$ and $L_{\text{Dice}}^{mask}$ denote the adaptive alignment losses between ${M}_{p}$ and $\hat{M}_{q}$.
@@ 4 Experiments
We evaluate our method on 3D referring datasets, such as ScanRefer [7], which is based on ScanNetv2 [10]. ScanRefer dataset consists of 1,513 intricately reconstructed 3D indoor scenes, following the official ScanNet splits. Additional experiments on SR3D/NR3D datasets and more qualitative results are detailed in the supplementary material.
Table 1: The 3D visual grounding results on ScanRefer, accuracy evaluated by IoU 0.25 and IoU 0.5. As depicted in the table, our method attains state-of-the-art (SOTA) performance in both two-stage and single-stage methods. Notably, in two-stage methods, our approach outperforms others in all performance metrics by a large margin.
@@ 4.2 Quantitative comparisons
In this subsection, we perform a quantitative comparison on ScanRefer, experiments show that our method achieves sota performance in both 3DREC and 3DRES tasks.
3DREC. We adopt Acc@0.25,Acc@0.5 as indicators to evaluate the 3DREC performance. As shown in Tab. 1, our method consistently exhibits superior performance over these prevailing methods in both two-stage and single-stage methods. Notably, in the realm of two-stage methods, our approach outperforms others in all performance metrics by a large margin (2.58% in Acc@0.25 and 3.27% in Acc@0.25). This evidence strongly suggests that the utilization of two separate branches facilitates the learning of task-specific information, leading to a substantial improvement in the performance of the 3DREC task through collaborative training between the branches. Furthermore, the remarkable results can be largely attributed to the well-designed Relative Superpoint Aggregation (RSA) module and Adaptive Soft Alignment (ASA) module, underscoring their effectiveness in the overall performance of the model. Nevertheless, in the single-stage approach, there is a decrease in performance under unique scenes. This could be attributed to the heightened recognition capability in complex scenes, potentially resulting in a lack of adaptability to simpler scenes.
3DRES. We utilize the evaluation metrics proposed earlier: mean IOU and Acc@kIOU. As demonstrated in Tab. 2, our method showcases a substantial performance improvement when compared to the previous method TGNN [17], as well as the latest approaches X-RefSeg3D [36] and 3D-STMN [40]. Notably, our method achieves impressive increases of 4.10% in Acc@0.25, 10.90% in Acc@0.5, and 5.22% in mIOU. Our model demonstrates superior performance, particularly in multiple scenes, with a huge improvement of 16.68% in Acc@0.5. This highlights the efficacy of aligning the two branches, enabling the model to distinguish and comprehend relationships in complex scenes.
@@ 4.3 Ablation Study
In the Relative Superpoint Aggregation (RSA) module, we try two different methods to fuse relative position vectors with point cloud features: “Plus” is shown in Equ. 2, “Mul” refers to the process of multiplying relative position vectors with point cloud features, followed by adding the fused features to the original point cloud features. As shown in Tab. 4, “Plus” surpasses “Mul”, especially in 3DREC metrics. This may be because the “Mul” fusion method induces substantial alterations to point features. As a result, the superpoint features obtained through the “Mul” fusion differ significantly from the original point cloud features. Consequently, during the alignment of the two modules, the outputs from the two branches do not align well, ultimately leading to suboptimal performance. The “Plus” fusion method not only incorporates the relative spatial information of superpoints but also endeavours to preserve the original features of the points to a significant extent.
@@
We performed an ablation experiment to investigate the impact of the Adaptive Soft Alignment (ASA) module and the Relative Superpoint Aggregation (RSA) module. The results, as depicted in Tab. 6, reveal that the basic MCLN, initially equipped with two branches for the 3DREC and 3DRES tasks and simply using ball query for Superpoint Aggregation, has already surpassed the performance of 3D-STMN [40] and EDA [41]. Notably, upon the integration of the ASA module, our model exhibits a substantial performance boost in all metrics. This suggests that our ASA module effectively facilitates the two branches in learning complementary information and enhances their coordination during training. Moreover, the integration of the ASA module with the RSA module results in a notable improvement in our model’s performance, particularly in the 3DRES task. This demonstrates that our RSA module contributes to a more precise aggregation of superpoint features, underscoring the efficacy of the RSA module. Additionally, the ASA module’s effectiveness in alignment is also substantiated.
@@
Figure 3: Qualitative results from 3D-STMN [40], EDA  [41] and our model. Our model exhibits superior positioning ability compared to 3D-STMN [40] and EDA  [41], along with a high level of consistency between the referred box and mask.
@@ A.1 More Performance Comparisons
SR3D/NR3D is derived from ScanNet, encompassing 83,572 automated, straightforward machine-generated descriptions in the case of SR3D and 41,503 descriptions in NR3D, resembling the human annotations employed in ScanRefer. The primary difference lies in the ScanRefer setup, where the task entails both object detection and matching. Conversely, SR3D/NR3D does not require this dual process. Tab. 7 presents the accuracy results on the SR3D and NR3D datasets, demonstrating our outstanding performance with $68.4\%$ and $59.8\%$ of Acc@0.25 in the 3DREC task. Especially in the NR3D dataset, our model overperforms the EDA [41] by a large margin of 6.7%, proving the effectiveness of our method. In particular, as depicted in Table 10, our approach consistently demonstrates exceptional performance in 3DRES tasks, achieving noteworthy mIOU of 49.84% and 46.09% in SR3D/NR3D dataset. These results underscore the universality of our method, emphasizing its superior predictive capabilities across both 3DREC and 3DRES tasks.

### v2

@@ Abstract
3D referring expression comprehension (3DREC) and segmentation (3DRES) have overlapping objectives, indicating their potential for collaboration. However, existing collaborative approaches predominantly depend on the results of one task to make predictions for the other, limiting effective collaboration. We argue that employing separate branches for 3DREC and 3DRES tasks enhances the model’s capacity to learn specific information for each task, enabling them to acquire complementary knowledge. Thus, we propose the MCLN framework, which includes independent branches for 3DREC and 3DRES tasks. This enables dedicated exploration of each task and effective coordination between the branches. Furthermore, to facilitate mutual reinforcement between these branches, we introduce a Relative Superpoint Aggregation (RSA) module and an Adaptive Soft Alignment (ASA) module. These modules significantly contribute to the precise alignment of prediction results from the two branches, directing the module to allocate increased attention to key positions. Comprehensive experimental evaluation demonstrates that our proposed method achieves state-of-the-art performance on both the 3DREC and 3DRES tasks, with an increase of 2.05% in Acc@0.5 for 3DREC and 3.96% in mIoU for 3DRES. Our code is available at https://github.com/qzp2018/MCLN.
@@
Figure 1: TGNN predicts the results of 3DREC and 3DRES tasks in a cascade manner. Our method introduces separate branches for 3DREC and 3DRES tasks and uses a parallel structure to facilitate collaborative training of them.
@@ 1 Introduction
Hence, it is intuitive to explore the feasibility of incorporating distinct branches dedicated to predicting 3DREC and 3DRES outcomes while training them collaboratively. To solve this issue, we use superpoints to facilitate the joint training of 3DREC and 3DRES tasks. The Relative Superpoint Aggregation (RSA) module is proposed to effectively exploit the spatial relationships between superpoints and their surrounding points during the feature aggregation process. Such an approach could potentially mitigate individual shortcomings through mutual compensation. However, merely training these branches together is insufficient, as the separate branches introduce the risk of prediction conflicts, which could jeopardize the overall performance. Therefore, ensuring alignment in the generated results from both branches is crucial. Furthermore, achieving alignment between the branches is vital in facilitating the exchange of bidirectional information between 3DREC and 3DRES branches. This collaborative learning approach empowers the branches to extract valuable information from one another, refining their positioning and contributing to a more effective performance.
@@
  • Our proposed method has achieved state-of-the-art performance in both 3DREC and 3DRES tasks, showcasing a notable improvement of 2.05% Acc@0.5 for 3DREC and a substantial increase of 3.96% mIoU for 3DRES.
@@ 2 Related Work
3D vision and language play pivotal roles in human comprehension of the world, and they serve as significant areas of research for advancing the development of machines with human-like capabilities. Consequently, noteworthy research endeavours in vision and language [42, 19, 20, 21, 31, 32, 30] have surged. Among these, 3D Visual Grounding (3DVG) [46, 44, 41, 17, 7, 5, 24, 9] stands out as a prime exemplar, focusing on the task of precisely localizing objects referred to in natural language descriptions. In a comprehensive scope, 3DVG comprises two fundamental tasks: 3D Referring Expression Comprehension (3DREC) [46, 44, 41, 7, 5, 45, 9] and 3D Referring Expression Segmentation (3DRES) [17, 40, 36]. For 3DREC, most of the existing works [7, 44, 45, 5] take a two-stage approach. Initially, candidate proposals are obtained using a pre-trained 3D object detector or ground truth information. Subsequently, visual and text features are extracted for multi-modal feature fusion, and then the most suitable proposal is selected based on the fused features. Recently, single-stage methods [41, 18] have also shown superior performance, presenting another promising paradigm. Conversely, although the 3DREC task has received considerable attention, the 3DRES task was initially introduced by TGNN [17], with limited exploration in subsequent years. Recently there has been a flurry of work to investigate this area. For example, 3D-STMN [40] employs superpoints for the 3DRES task, greatly improving performance. However, the collaborative work between 3DREC and 3DRES is still very few, indicating significant untapped potential.
@@ 2.2 Multi-task Learning
Multi-task learning (MTL) [28, 6, 22, 34, 16] serves as a valuable strategy for enhancing task performance, effectively leveraging training data while facilitating the exchange of information among related tasks. MTL has found extensive applications across a diverse range of computer vision tasks. In the 2D vision, many characteristics are shared between detection and segmentation, which have prompted several works [15, 4, 12, 24] to address both tasks within a unified framework. Similarly, REC and RES can also promote each other in joint training [28, 16, 22]. However, due to the limited exploration of 3DRES in recent years, collaborative initiatives between 3DRES and 3DREC have mainly remained uncharted.
TGNN [17] conducts the first attempt to employ a single model to tackle both 3DREC and 3DRES tasks. However, it doesn’t set up a separate branch for the 3DREC task, and this design exhibits several limitations. Specifically, it ties the performance ceiling for 3DREC solely to 3DRES, overlooking the distinctions between these tasks. Similarly, 3DRefTR [24] also adopts the cascade structure to tackle two tasks, severely restricting the interaction between the 3DREC and 3DRES tasks and leading to sub-optimal performance. To deal with the above problems, we designed two separate branches for 3DREC and 3DRES and a co-alignment module for joint training.
@@ 3 Method
Figure 2: An overview of our proposed network. Our proposed network comprises a cross-modal encoder, separate decoders for 3DREC and 3DRES, a “Relative Superpoint Aggregation” (RSA) module for superpoint feature generation, and an “Adaptive Soft Alignment” (ASA) module for joint training.
@@ 3.3 Adaptive Soft Alignment
The query mask $\hat{M}_{q}$ acts as an indirect representation of the predicted bounding box. It serves as a connecting link, facilitating the alignment between the bounding box from the 3DREC branch and the mask from the 3DRES branch. This soft alignment scheme enhances the coherence between the two branches. Through the alignment of superpoint masks ${M}_{p}$ and $\hat{M}_{q}$, the indirect preservation of spatial coherence between the outcomes of the 3DREC and 3DRES tasks is effectively achieved. Usually, a Focal loss [25] and a Dice loss [33] will be applied to conduct the alignment between masks.
@@ 3.4 Overall Loss
$L_{\text{res}}=\beta_{1}(L_{\text{Focal}}^{1}+L_{\text{Focal}}^{2}+L_{\text{Focal}}^{fin}+L_{\text{Focal}}^{point})+\beta_{2}(L_{\text{Dice}}^{1}+L_{\text{Dice}}^{2}+L_{\text{Dice}}^{fin}+L_{\text{Dice}}^{mask}),$ (18)
where $L_{\text{Focal}}^{1},L_{\text{Dice}}^{1}$ denotes the losses associated with the superpoint mask ${M}_{f}$ and the ground truth. Similarly, $L_{\text{Focal}}^{2},L_{\text{Dice}}^{2}$ represents the losses for the superpoint mask ${M}_{q}$. Additionally, $L_{\text{Focal}}^{fin},L_{\text{Dice}}^{fin}$ correspond to the losses for the final predicted superpoint mask ${M}_{f}$. Finally, $L_{\text{Focal}}^{point}$ and $L_{\text{Dice}}^{mask}$ denote the adaptive alignment losses between ${M}_{p}$ and $\hat{M}_{q}$.
@@ 4 Experiments
We evaluate our method on 3D referring datasets, such as ScanRefer [7]. ScanRefer dataset consists of 1,513 intricately reconstructed 3D indoor scenes, following the official ScanNet [10] splits. Additional experiments on SR3D/NR3D datasets and more qualitative results are detailed in the Appendix.
Table 1: The 3D visual grounding results on ScanRefer, accuracy evaluated by IoU 0.25 and IoU 0.5. As depicted in the table, our method attains state-of-the-art (SOTA) performance in both two-stage and single-stage methods.
@@ 4.2 Quantitative comparisons
In this subsection, experiments on ScanRefer show that our method achieves SOTA performance in both 3DREC and 3DRES tasks.
3DREC. We adopt Acc@0.25,Acc@0.5 as indicators to evaluate the 3DREC performance. As shown in Tab. 1, our method consistently exhibits superior performance over these prevailing methods in both two-stage and single-stage methods. Notably, in the realm of two-stage methods, our approach outperforms others in all performance metrics by a large margin (2.05% in Acc@0.5). This evidence strongly suggests that the utilization of two separate branches facilitates the learning of task-specific information, leading to a substantial improvement in the performance of the 3DREC task through collaborative training between the branches. Furthermore, the remarkable results can be largely attributed to the well-designed Relative Superpoint Aggregation (RSA) module and Adaptive Soft Alignment (ASA) module, underscoring their effectiveness in the overall performance of the model. Nevertheless, in the single-stage approach, there is a decrease in performance under unique scenes. This could be attributed to the heightened recognition capability in complex scenes, potentially resulting in a lack of adaptability to simpler scenes.
3DRES. We utilize the evaluation metrics proposed earlier: mean IoU and Acc@kIoU. As demonstrated in Tab. 2, our method showcases a substantial performance improvement when compared to the previous method TGNN [17], as well as the latest approaches [36, 40, 24]. Notably, our method achieves impressive increases of 4.63% in Acc@0.5, and 3.96% in mIoU. This highlights the efficacy of aligning the two branches, enabling the model to distinguish and comprehend relationships in complex scenes.
@@ 4.3 Ablation Study
In the Relative Superpoint Aggregation (RSA) module, we try two different methods to fuse relative position vectors with point cloud features: “Plus” is shown in Equ. 2, “Mul” refers to the process of multiplying relative position vectors with point cloud features, followed by adding the fused features to the original point cloud features. As shown in Tab. 4, “Plus” surpasses “Mul”, especially in 3DREC metrics. This may be because the “Mul” fusion method induces substantial alterations to point features. As a result, the superpoint features obtained through the “Mul” fusion differ significantly from the original point cloud features. Consequently, during the alignment of the two modules, the outputs from the two branches don’t align well, leading to suboptimal performance. The “Plus” fusion method not only incorporates the relative spatial information of superpoints but also endeavours to preserve the original features of the points to a significant extent.
@@
We performed an ablation experiment to investigate the impact of the Adaptive Soft Alignment (ASA) module and the Relative Superpoint Aggregation (RSA) module. The results, as depicted in Tab. 6, reveal that the basic MCLN, initially equipped with two branches for the 3DREC and 3DRES tasks and simply using ball query for Superpoint Aggregation, has already surpassed the performance of 3D-STMN [40] and EDA [41]. Notably, upon the integration of the ASA module, our model exhibits a substantial performance boost in all metrics. This suggests that our ASA module effectively facilitates the two branches in learning complementary information and enhances their coordination during training. Moreover, the ASA module results in a notable improvement in our model’s performance, particularly in the 3DRES task. This demonstrates that our RSA module contributes to a more precise aggregation of superpoint features, underscoring the efficacy of the RSA module. Additionally, the ASA module’s effectiveness in alignment is also substantiated.
@@
Figure 3: Qualitative results from 3D-STMN [40], EDA  [41] and our model.
@@ A.1 More Performance Comparisons
SR3D/NR3D is derived from ScanNet, encompassing 83,572 automated, straightforward machine-generated descriptions in the case of SR3D and 41,503 descriptions in NR3D, resembling the human annotations employed in ScanRefer. The primary difference lies in the ScanRefer setup, where the task entails both object detection and matching. Conversely, SR3D/NR3D does not require this dual process. Tab. 7 presents the accuracy results on the SR3D and NR3D datasets, demonstrating our outstanding performance with $68.4\%$ and $59.8\%$ of Acc@0.25 in the 3DREC task. Especially in the NR3D dataset, our model overperforms the 3DRefTR [24] by a large margin of 7.2%, proving the effectiveness of our method. In particular, as depicted in Table 10, our approach consistently demonstrates exceptional performance in 3DRES tasks, achieving noteworthy mIOU of 49.84% and 46.09% in SR3D/NR3D dataset. These results underscore the universality of our method, emphasizing its superior predictive capabilities across both 3DREC and 3DRES tasks.